OpenStudy (anonymous):

-2(x+1)-3 is greater than or equal to 11

OpenStudy (jdoe0001):

ok, thanks for sharing that :)

OpenStudy (jdoe0001):

$$ -2(x+1)-3 \ge 11\\ \text{just work it as you'd }\\ -2(x+1)-3 = 11\\ \text{just using the } \ge \ sign $$

OpenStudy (irishboy123):

start by expanding the left side of the equation. it currently looks so messy, but that is the challenge that your examiners are concerned with: -2(x+1)-3= -2x -2 -3 = -2x -5 OK we have sorted out the left hand side, so back to the main game: -2x -5 >= 11 add 5 to each side maintaining the equality/ inequality -2x-5+5 >= 11+5 -2x >= 16 divide EACH side by 2, maintaining the equality/ inequality -2x/2 >= 16/2 -x>=8 this the the annoying bit. to convert -x to x (because your teacher wants to know the value of x), you need to multiply EACH side by -1. that would never change an = sign, but this is an >= sign. when you multiply by -1, > becomes <, and < becomes >. Equally, => becomes =<, and =< becomes =>. Use a number line to see why. so, if we multiply by -1, we get: x <= -8 does that make sense?